OpenStudy (anonymous):

Through (3,2);perpendicular to 2x+y=1

OpenStudy (anonymous):

Get that equation in slope-intercept form first, then take the slope you have, flip it and change the sign. Keep the same y-intercept and that's your equation

OpenStudy (aum):

\[ 2x + y = 1 \\ y = -2x + 1 \\ \]Slope = -2. Slope of perpendicular line is 1/2. y = mx + b y = 1/2x + b. It passes through the point (3,2). When x = 3, y = 2: 2 = 1/2 * 3 + b b = 2 - 3/2 = 1/2 y = 1/2x + 1/2 is the equation of the perpendicular line.